ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 101, 1997

Abramovich, D. Lang maps and Harris’s conjecture  85

Brudnyi, Y. and A. Shteinberg Calderón constants of finitedimensional couples

289

Cutillas Ripoll, P. and J. M. Verde Ramírez Weierstrass factorizations in compact Riemann surfaces

205

Drasin, D. On a method of Holopainen and Rickman  73

Dugas, M. and R. Göbel Endomorphism rings of B_{2}-groups of infinite rank

141

El Kacimi Alaoui, A. et B. Gmira Stabilité du caractère kählérien transverse

323

Gemira, B. See A. El Kacimi Alaoui

Göbel, R. See M. Dugas

Guralnick, R. and D. Wan Bounds for fixed point free elements in a transitive group and applications to curves over finite fields

255

Kalai, G. and J. Matous̆ek Guarding galleries where every point sees a large area

125

Marrero, J. C. and E. Padrón Compact generalized Hopf and cosymplectic solvmanifolds and the Heisenberg group H(n, 1)

189

Matous̆ek, J. See G. Kalai

Padrón, E. See J. C. Marrero

Rudelson, M. Contact points of convex bodies  93
Schneider, D. Théorèmes ergodiques perturbes  157
Shlapentokh, A. The logic of pseudo-S-integers  229

Shteinberg, A. See Y. Brudnyi

Verde Ramírez, J. M. See P. Cutillas Ripoll

Vo, S. C. The spin L-function on the symplectic group GSp(6)  1

Wan, D. See R. Guralnick

Zampieri, G. Hypersurfaces through higher-codimensional submanifolds of ℂ^{n} with preserved Levi-kernel 179